OpenStudy (anonymous):

The length of the curve y=x^3 from x=0 to x-2 is given by

OpenStudy (amistre64):

\[\int_{a}^{b}ds\] \[ds = \sqrt{(\frac{dx}{dt})^2+(\frac{dy}{dt})^2}~ dt\] let x=t and y=t^3